OpenStudy (anonymous):

Find the slope of the line that passes through the points (2, 1) and (-4, -5).

OpenStudy (anonymous):

(-6,-6)

OpenStudy (anonymous):

That isn't the answer...

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Ah, okay! Thanks. Here's one more: (0, 0) and (-2, -3)

OpenStudy (anonymous):

I'm gonna laugh if it's 2/3

OpenStudy (abb0t):

\[\frac{Y_2 - Y_1 }{ X_2 - X_1 } = m \] where m is your slope.

OpenStudy (anonymous):

m=3/2

OpenStudy (anonymous):

Thanks. And abb0t, thanks for reminding me of the slope equation. But I have to give best response to the one who helped out first.